OpenStudy (anonymous):

Put the following radical expression into simplified form. @kx2bay

OpenStudy (anonymous):

\[\sqrt{\frac{ 7 }{ 5 }}\]

OpenStudy (anonymous):

can it get simpler that that? what do they want...

OpenStudy (anonymous):

I have no idea lol

OpenStudy (anonymous):

1.1832 may be...

OpenStudy (anonymous):

Ok, on an example in the book, we have to use prop 2, to separate them: \[\sqrt{\frac{ 7 }{ 5 }}=\frac{ \sqrt{7} }{ \sqrt{5} }\] Then multiply both #s by \[\sqrt{5}\] \[\frac{ \sqrt{7} }{ \sqrt{5} }*\frac{ \sqrt{5} }{ \sqrt{5} }\]

OpenStudy (anonymous):

so it will become \[\sqrt{35} / 5\]

OpenStudy (anonymous):

yup ! lol

OpenStudy (anonymous):

may be you are right cookii and that's what they're after, well done, u little genius :)

OpenStudy (anonymous):

Let me check the answer first lol

OpenStudy (anonymous):

yes yes good idea...

OpenStudy (anonymous):

Correct lol

OpenStudy (anonymous):

told u...you becoming a smartie now

OpenStudy (anonymous):

still hate math though lol

OpenStudy (anonymous):

well you have to hate it to love it, that's math for u :)

OpenStudy (anonymous):

lol